OpenStudy (anonymous):

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $ 123.00 for 3 days and 300 miles, while Mary was charged $ 216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage?

OpenStudy (dumbcow):

x is daily fee, y is mileage fee 3x +300y = 123 5x +600y = 216 use elimination multiply top by -2 -6x -600y = -246 5x +600y = 216 _____________________ -x +0 = -30 x = 30 find y by replacing x with 30 3(30) +300y = 123 300y = 123-90 = 33 y = 33/300 = 11/100

OpenStudy (dumbcow):

11/100 = 0.11 so 11 cents per mile